OpenStudy (anonymous):

Divide. (-6m7 + 20m6) ÷ 2m4

OpenStudy (anonymous):

\[\frac{20 m^6-6 m^7}{2 m^4}=\frac{1}{2 m^4} \left(20 m^6-6 m^7\right)=10 m^2-3 m^3 \]